Consider the following reversible reaction occurring in a 1.00 L
enclosed vessel:
(NH4)2SO4(s) to 2NH3(g) + H2SO4(g)
Given that the value of the equilibrium constant, Kc, of this reaction is
2.57 × 10-5
, calculate the molar concentrations of NH3(g) and H2SO4(g)
at equilibrium. No NH3(g) and H2SO4(g) were present initially.
1
Expert's answer
2020-08-13T07:48:32-0400

Let concentration of H2SO4 IS X.


CONCENTRATION OF NH3 IS 2X.


So , Equillibrium Constant ( Kc ) . = (2X)2 . (X)1 . = 4(X)3 .


4X3 = 2.57 * 10-5 ..


X = 2.57 * 10-5 / 4 .


X = 6 * 10-6 .


Concentration of H2SO4 . = X = 6 * 10-6 .


Concentration of NH3 = 2X = 12 * 10-6 .